OpenStudy (anonymous):

simplify (42xy)/(54x)

OpenStudy (anonymous):

simplify the coeficent and cancell out your x varible

OpenStudy (anonymous):

how to 42 x 54

OpenStudy (anonymous):

(42xy)/(54x)

OpenStudy (anonymous):

are you dividing this ? because thats what it looks like with the "/" sign in the middle

OpenStudy (anonymous):

(42xy)/(54x) =7y/9

OpenStudy (anonymous):

\[\frac{7*3*2 *x *y}{3*3*3*2*x}=\frac{7 y}{9} \]

OpenStudy (anonymous):

Sorry. Two editions was not intended.